# Pinpointing Important Genetic Variants via A Feature-level Filter and Group Knockoffs

### v2

@@ 2.4 Rigorous Feature Filters under Group Knockoffs
[…]
@@
Based on such an observation, we utilize theoretical results of Katsevich and Sabatti, (2019) that for any $t>0$,
@@
[…]
@@
The proof of Theorem 2 is provided in Section B of the Appendix. Similar to the discussion in Sesia et al., (2020), the correction factor $1.93$ is required in the proof for technical reasons, while empirical evidence suggests that it may be practically unnecessary. To solve the optimization problem (10), we develop the feature-versus-group (FVG) filter (Algorithm 2) in an analogous manner to the GGM knockoff filter (Li and Maathuis,, 2021).
We also provide other variations for FDR control with details in Section C of the Appendix.
@@
In the FVG filter (Algorithm 2), the budgets $v_{l}$’s can be computed adaptively to the prior belief of the number of important features within each groups. Here, we provide two strategies of specifying $v_{l}$’s.
@@
3 Simulated Experiments
@@
To evaluate the proposed FVG filter in both FDR control and identifying important features, we conduct extensive experiments under various scenarios. Specifically, throughout simulated experiments, we generate group knockoffs under the maximum entropy (ME) construction of Chu et al., (2023), compute feature statistics as $W_{j}=\text{sign}\left(T_{j}-\widetilde{T}_{j}\right)\cdot\max\left(T_{j},\widetilde{T}_{j}\right)$ with the joint lasso importance scores.
3.1 Simulated Experiments with Synthetic Data
To examine how the FVG filter performs under different sample sizes, we simulate 1000 datasets with a fixed number of features $p=250$ for sample sizes $n=500,1000$ and $2000$. Specifically, we partition these $250$ features into $K=50$ groups of equal size $B_{1}=\{X_{1},\ldots,X_{5}\},\ldots,B_{50}=\{X_{246},\ldots,X_{250}\}$. Based on such a partition, we generate for each simulated dataset $n$ feature vectors $\textbf{x}_{1},\ldots,\textbf{x}_{n}$ from the multivariate normal distribution,
$\text{MVN}\left\{\mathbf{0},\begin{pmatrix}0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ 0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ \vdots&\vdots&\ddots&\vdots\\ 0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ \end{pmatrix}\right\}.$ (12)
Given features $\textbf{x}_{1},\ldots,\textbf{x}_{n}$, responses $y_{1},\ldots,y_{n}$ are simulated from the linear model
$Y=\sum_{k=1}^{K}\textbf{X}_{B_{k}}^{\mathsf{T}}\boldsymbol{\beta}_{B_{k}}+e,\quad e\sim\text{N}(0,4^{2}),$ (13)
where
$\boldsymbol{\beta}_{B_{k}}=\begin{cases}(\textbf{1}_{2}^{\mathsf{T}},-0.2\cdot\textbf{1}_{3}^{\mathsf{T}})^{\mathsf{T}},&k=1,2,\\ (2.5,-1.2\cdot\textbf{1}_{2}^{\mathsf{T}},\textbf{0}_{2}^{\mathsf{T}})^{\mathsf{T}},&k=3,4,\\ (0.3\cdot\textbf{1}_{2}^{\mathsf{T}},\textbf{0}_{3}^{\mathsf{T}})^{\mathsf{T}},&k=5,\ldots,10\\ \textbf{0}_{5},&k=11,\ldots,50,\\ \end{cases}$ (14)
Under (12)-(14), we have $H_{j}^{(\text{fg})}$’s are false for $j=51,\ldots,250$ and there are 28 important features. Empirical FDR and power of the FVG filter in identifying important features under different sample sizes are visualized in Figure 1. For comparison, we also evaluate the performance of model-X knockoff filter (Candès et al.,, 2018), group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019) in FDR control and identifying important features.
@@ 3.1 Simulated Experiments with Synthetic Data
Figure 1: Empirical FDR and power of FVG filter with respect to the target FDR level ($\alpha$) over 1000 simulated datasets of different sample sizes in comparison to model-X knockoff filter (Candès et al.,, 2018), group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019).
It is clear that the FVG filter manages to control $\text{FDR}^{(\text{fg})}$ under any target level $\alpha$. For a fixed target level $\alpha$, we find the empirical power increases to $1$ as $n$ grows, suggesting that the FVG filter is consistent in identifying important features. Although existing methods can also control FDR and have growing power, they suffer different types of performance deterioration in identifying important features. On one hand, group knockoff filter suffers power loss when the target FDR level is lower than $0.10$ due to the threshold issue mentioned in Gimenez and Zou, (2019). Even with power improvement as the target FDR level grows from $0.10$ and $0.20$, the informativeness of rejection sets obtained by group knockoff filter is poor as all important groups ($B_{1},\ldots,B_{10}$) are rejected as a whole without distinguishing which features are more important. As a comparison, the average size of nonempty catching sets $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ ($k=1,\ldots,50$) obtained by FVG filter is less than $2$ under the target level $\alpha=0.1$. On the other hand, although the multilayer knockoff filter has power improvement over model-X knockoff filter by focusing important feature selection on those identified feature groups, there remains a significant power gap between FVG filter and the multilayer knockoff filter when the target FDR is large.
3.2 Simulated Experiments with Real-world Genetic Data
To mimic the dependency structure among features in real-world genetic analysis and investigate how the proposed method performs in real applications, we simulate 500 datasets based on $6095$ variants in the APOE/APOC region (chromosome 19, positions 44909011 $\sim$ 45912650) from the whole-genome sequencing data (NG00067.v5) of the Alzheimer’s Disease Sequencing Project (Leung et al.,, 2019). Specifically, we restrict our study to $6952$ individuals with estimated European ancestry rate [^3] Estimated by SNPWeights v2.1 (Chen et al.,, 2013) using reference populations from the 1000 Genomes Consortium (The 1000 Genomes Project Consortium,, 2015). greater than 80% and $p=1000$ variants $X_{1},\ldots,X_{p}$ whose minor allele frequency (MAF) is larger than $1\%$ and pairwise correlations are in $[-0.95,0.95]$ to avoid the existence of statistically indistinguishable variants with nearly-perfect correlations. Among these variants, we compute the correlation $\text{cor}(X_{i},X_{j})$ between any pair of variants $(X_{i},X_{j})$ over all 6952 individuals and construct variant groups ${B}_{1},\ldots,{B}_{K}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(X_{i},X_{j})|)_{p\times p}$. By randomly sampling $n=5000$ individuals without replacement, each simulated dataset is obtained by collecting the sampled individuals’ variants and generating responses $y_{1},\ldots,y_{n}$ from the linear model
$Y=\sum_{j=1}^{p}X_{j}{\beta}_{j}+e,\quad e\sim\text{N}(0,4),$ (15)
where only $k=10$ randomly selected coefficients ${\beta}_{j}$’s are nonzero and follow $N(0,0.225)$. As a result, important features refer to variants $X_{j}$’s whose corresponding coefficients $\beta_{j}$’s are nonzero and important groups are those containing at least one important feature. Specifically, most important groups only contain one important feature.
Figure 2: Empirical performance of the model-X knockoff filter (Candès et al.,, 2018), the group knockoff filter (Dai and Barber,, 2016), the multilayer knockoff filter (Katsevich and Sabatti,, 2019) and our FVG filter over 500 simulated data: (a) empirical FDR and power with respect to the target FDR level; (b) average size of catching sets; (c) average purity of catching sets.
To illustrate the empirical benefit of the FVG filter, we apply it to find important features in these 500 simulated datasets with comparison to the model-X knockoff filter (Candès et al.,, 2018), the group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019). Empirical FDR and power of all methods are presented in Figure 2 (a). For fair comparison, we use the joint lasso feature importance score when implementing both existing methods and our algorithm. It is found that our FVG filter can empirically control FDR. Compared to the model-X knockoff filter which can only distinguish around $50\%$ of important features when the target FDR level is relaxed to $20\%$, our FVG filter, the multilayer knockoff filter and the group knockoff filter are able to identify about $70\%\sim 80\%$ of important features (or feature groups) when the target FDR level is $10\%$. However, the inference results given by the multilayer knockoff filter and the group knockoff filter are far from being informative and precise. To illustrate this, we compute catching set $\mathcal{R}_{k}$ of each variant group from the rejection set $\mathcal{R}$, where
  • $\mathcal{R}_{k}=B_{k}$ if $H^{\text{(gg)}}_{k}$ is rejected by the group knockoff filter;
  • $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ if the rejection set $\mathcal{R}$ is obtained from the model-X knockoff filter or our FVG filter.
  • For each group $B_{k}$, we let
    • $\mathcal{R}_{k}=B_{k}$ if $H^{\text{(gg)}}_{k}$ but all $H^{\text{(ff)}}_{j}$’s ($j\in B_{k}$) is rejected by the multilayer knockoff filter;
    • $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ if $H^{\text{(gg)}}_{k}$ and some $H^{\text{(ff)}}_{j}$’s ($j\in B_{k}$) are rejected by the multilayer knockoff filter.
We then calculate the average size and purity (minimum absolute correlation) of catching sets to measure the informativeness and precision of $\mathcal{R}$ given by different filters in Figure 2 (b)-(c). Here, it is desired that sizes of catching sets are smaller as more unimportant variants are excluded. Even when there remain multiple features in catching sets, high purity is desirable to exclude unimportant variants weakly correlated to the important ones (or poor-quality proxy variants). In contrast to the group knockoff filter whose average size and purity of catching sets are almost $6$ and smaller than $20\%$ respectively (which means that catching sets obtained by the group knockoff filter contain a lot of poor-quality proxy variants), the precision of our FVG filter in pinpointing true important variants is close to the model-X knockoff filter and is higher than the multilayer knockoff filter. With an average size close to $1$ and average purity greater than $80\%$, our FVG filter can directly obtain informative catching sets containing only one or two features in most cases. In summary, our FVG filter manage to simultaneously achieve high power and high informativeness in identifying important features.

### v3

@@ 2.4 FVG Filter
$\max|\mathcal{R}^{(\text{fg})}|,\quad\text{s.t. }$ $\widehat{\text{FDP}}^{(\text{fg})}=\sum_{l}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\times\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\leq\alpha,$
@@
Based on such an observation, we utilize theoretical results of Katsevich2019 that for any $t>0$,
@@
$\max|\mathcal{R}^{(\text{fg})}|,\quad\text{s.t. }$ $\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\times\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\leq\frac{v_{l}\alpha}{1.93},\quad l=1,2,\ldots,$ (10)
@@
The proof of Theorem 2 is provided in Appendix C. Observing that the optimization problem (10) is analogous to the ones in (19) and (21) of Li2021 that
  • $\star$
    feature statistics $W_{j}$’s within the each row of Table 2 satisfy the coin-flipping property while $W_{j}$’s for $j=1,\ldots,p$ do not (which is analogous to the $p\times p$ feature statistics matrix in Section 3.1 of Li2021) and;
  • $\star$
    the contribution of each row of Table 2 to $\widehat{\text{FDP}}^{(\text{fg})}$, measured by
    $\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\times\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]},$
    is required to be constrained under the row-specific level $\frac{v_{l}\alpha}{1.93}$ with $\sum_{l}\frac{v_{l}\alpha}{1.93}=\frac{\alpha}{1.93}$,
we develop the feature-versus-group (FVG) filter (Algorithm 2), whose steps 4-10 adopt the threshold calculation strategy in Algorithm 2 of Li2021 to compute the optimal thresholds $t^{(l)}$’s. Similar to the discussion in supplementary note 4 of Sesia2020, the correction factor $1.93$ is required in the proof for technical reasons, while empirical evidence suggests that it may be practically unnecessary. Thus, for all experiments in Sections 34, we implement the FVG filter without the correction factor $1.93$. For empirical reference, we compare the naive filter (Algorithm 1), the FVG filter with and without the $1.93$ factor via experiments in Appendix D.
@@
We also provide other variations with details in Appendix E.
@@
In the FVG filter (Algorithm 2), the budgets $v_{l}$’s can be computed adaptively to the prior belief of the number of important features within each group. Here, we provide two strategies of specifying $v_{l}$’s.
@@
Unlike the FDR control, asymptotic power of our FVG filter can’t be rigorously proved. The main reason is that our FVG filter is a model-free approach without rigid parametric assumption on the conditional distribution $Y|\textbf{X}$ and its power relies on the consistency between $Y|\textbf{X}$ and the feature importance score. However, when a linear model is correctly used, our FVG filter can achieve asymptotic power $1$. For example, following the proof of Zou2006, the feature importance scores $W_{j}$’s will be positive with asymptotic probability $1$ for those features with nonzero contribution to $Y$, making their asymptotic probability to be included in the selection set $\mathcal{R}^{(\text{fg})}$ as 1 under mild conditions.
3 Simulated Experiments with Real-world Genetic Data
To evaluate the proposed FVG filter in controlling FDR and identifying important features with comparison to existing approaches, we conduct extensive simulated experiments.
3.1 Data Generation
We simulate datasets based on a real-world genetic data to mimic the dependency structure among features in real-world genetic analysis and investigate how the proposed method performs in real applications. Specifically, we simulate datasets based on $6095$ variants in the APOE/APOC region (chromosome 19, positions 44909011 $\sim$ 45912650) from the whole-genome sequencing data (NG00067.v5) of the Alzheimer’s Disease Sequencing Project (Leung2019). In our study, we restrict our focus to $6952$ individuals with estimated European ancestry rate [^2] Estimated by SNPWeights v2.1 (Chen2013) using reference populations from the 1000 Genomes Consortium (10002015). greater than 80% and $p=1157$ variants $X_{1},\ldots,X_{p}$ whose minor allele frequency (MAF) is larger than $1\%$ and pairwise correlations are in [0.95,0.95][-0.95,0.95] to avoid the existence of statistically indistinguishable variants with nearly-perfect correlations. By randomly sampling $n$ individuals without replacement, each simulated dataset is obtained by collecting the sampled individuals’ variants and generating responses $y_{1},\ldots,y_{n}$ from the linear model
$Y=\sum_{j=1}^{p}X_{j}{\beta}_{j}+e,\quad e\sim\text{N}(0,4),$ (12)
where only $k=15$ randomly selected coefficients ${\beta}_{j}$’s are nonzero and follow $N(0,0.6^{2})$. In our experiments, we simulate 1000 datasets for each possible sample size $n=500,1000$ and $2000$. To circumvent the threshold phenomenon mentioned in Gimenez2019b, we implement the multiple knockoff counterpart described in Appendix E.1.
3.2 Details of Implementation and Evaluation Metrics
Among these $1157$ variants, we compute the correlation $\text{cor}(X_{i},X_{j})$ between each pair of variants $(X_{i},X_{j})$ over all 6952 individuals and construct $345$ variant groups ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(X_{i},X_{j})|)_{p\times p}$. We define important features as variants $X_{j}$’s whose corresponding coefficients $\beta_{j}$’s are nonzero and important groups as those containing at least one important feature. Specifically, most important groups only contain one important feature. In summary, we have
  • $H_{j}^{(\text{ff})}$’s are false for 15 important features;
  • $H_{k}^{(\text{gg})}$’s are false for groups containing important features;
  • $H_{j}^{(\text{fg})}$’s are false for features in those groups with false $H_{k}^{(\text{gg})}$’s.
@@ 3.2 Details of Implementation and Evaluation Metrics
Based on groups ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ and the correlation matrix $\boldsymbol{\Sigma}=(\text{cor}(X_{i},X_{j}))_{1157\times 1157}$, we generate second-order group knockoffs under the maximum entropy (ME) construction using Algorithm 2 of Chu2024 and compute feature statistics as $W_{j}=\text{sign}\left(T_{j}-\widetilde{T}_{j}\right)\cdot\max\left(T_{j},\widetilde{T}_{j}\right)$ with the joint lasso importance scores in the implementation of our FVG filter.
To evaluate the power and precision of inference results produced by our FVG filter, we derive catching sets of signals based on the selection set $\mathcal{R}_{\text{FVG}}$ as
$\mathcal{CC}_{\text{FVG}}=\{\mathcal{S}_{k,\text{FVG}}={B}^{(2)}_{k}\cap\mathcal{R}_{\text{FVG}}|k=1,\ldots,345\}.$
For comparison, we also implement several existing methods as follows.
  • Model-X knockoff filter: Based on the correlation matrix $\boldsymbol{\Sigma}$, we generate second-order knockoffs at the variant level under the maximum entropy (ME) construction and use the lasso coefficient difference statistics (Candes2018) in the implementation of model-X knockoff filter. Based on the selection set $\mathcal{R}_{\text{model-X}}$, we derive catching sets as
    $\mathcal{CC}_{\text{model-X}}=\{\mathcal{S}_{j,\text{model-X}}=\{j\}|j\in\mathcal{R}_{\text{model-X}}\}.$
  • Group knockoff filter: We implement the group knockoff filter based on the same ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ and the same second-order group knockoffs as our FVG filter. Based on the selection set $\mathcal{R}_{\text{group}}$, we derive catching sets as
    $\mathcal{CC}_{\text{group}}=\{\mathcal{S}_{k,\text{group}}={B}^{(2)}_{k}|k\in\mathcal{R}_{\text{group}}\}.$
  • KnockoffZoom and KeLP: For these methods which identify signals at multiple layers of different resolutions, we define three resolution layers as (a) layer 0: the level of variants; (b) layer 1: the level of variant groups ${B}^{(1)}_{1},\ldots,{B}^{(1)}_{650}$ obtained by the hierarchical clustering (average linkage with cutoff value $0.25$); (c) layer 2: the level of variant groups ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ which are also used in the FVG filter. Here, layer 2 corresponds to the same variant group level as the group knockoff filter while layer 1 is an intermediate layers between layer 2 and the level of variants. Based on such a hierarchical structure of variant groups that each variant group at high-resolution levels is nested under one group at low-resolution levels, we implement both KnockoffZoom and KeLP at layers 0 and $2$ (2 layers) and layers 0, 11 and $2$ (3 layers). Thus, both KnockoffZoom and KeLP are tackling the same task as our FVG filter to refine those important variant groups in layer 2. For each layer, we generate second-order group knockoffs under the maximum entropy (ME) construction using Algorithm 2 of Chu2024. After implementing either KnockoffZoom or KeLP, we obtain catching sets as the most specific discoveries among all discoveries in selection sets (same as Figure 1 of Gablenz2024) as detailed in Appendix F. Specifically, to avoid obtaining empty catching sets, we implement KeLP in our article with carefully chosen tuning parameters as detailed in Appendix G.
In the following section, we evaluate the empirical FDR of our FVG filter via (4), empirical power of all approaches as the proportion of important features (with nonzero coefficients) included in at least one catching sets. We also measure the informativeness of different approaches by measuring the average size and purity [^3] Purity refers to the minimum absolute correlation within the catching set. of catching sets.
3.3 Performance Comparison of the FVG Filter with Existing Competitors
Refer to captionRefer to caption
@@ 3.3 Performance Comparison of the FVG Filter with Existing Competitors
Figure 2: Empirical FDR and power of FVG filter with respect to the target FDR level ($\alpha$) over 1000 simulated genetic datasets of different sample sizes in comparison to model-X knockoff filter (Candes2018), group knockoff filter (Dai2016), the KnockoffZoom (Sesia2020) and the KeLP (Gablenz2024).
Refer to captionRefer to caption
(a) $n=500$.
Refer to captionRefer to caption
(b) $n=1000$.
Refer to captionRefer to caption
(c) $n=2000$.
Figure 3: Average sizes and purity of catching sets obtained FVG filter over 1000 simulated genetic datasets of different sample sizes in comparison to model-X knockoff filter (Candes2018), group knockoff filter (Dai2016), the KnockoffZoom (Sesia2020) and the KeLP (Gablenz2024).
Over 1000 simulated datasets of different sample sizes, the empirical FDR of our FVG filter and its power in comparison with all existing competitors under different target FDR levels are reported in Figure 2. We also present the average sizes and purity of catching sets produced by different methods in Figure 3. It is clear from Figure 2 that the FVG filter controls $\text{FDR}^{(\text{fg})}$ at any target level $\alpha$.
Existing methods suffer different types of performance deterioration in identifying important features. On one hand, the model-X knockoff filter suffers great power loss in Figure 2 because high correlations of variants in the same group lower signal-to-noise ratio of false $H^{\text{(ff)}}_{j}$’s. On the other hand, although group knockoff filter manages to achieve higher power in Figure 2, such an advantage comes from the sacrifice of informativeness of rejection sets. This can be seen from catching sets with average size of around $6$ and average purity of around $0.6$ in Figure 3 because variant groups are rejected as a whole without distinguishing which variants are more important. Such a poor precision of group knockoff filter can be somehow relieved by both KnockoffZoom and KeLP, which however still suffer lack of precision. Specifically, KnockoffZoom (2 layers) manages to maintain similar power as the group knockoff filter as shown in Figure 2. By decreasing the average size from $6$ to $3$ and increasing average purity from $0.6$ to $0.85$ when $n=2000$, the improvement of precision provided by KnockoffZoom is still limited. Adding the intermediate layer (layer 1) in KnockoffZoom only slightly improves precision of catching sets (average size of around $2.5$ and average purity of around $0.89$) without any power gain. KeLP, on the other hand, improves the precision of catching sets much greater than KnockoffZoom. With the same 2 layers in analysis, KeLP (2 layers) returns catching sets of average sizes of around $2.3$ and average purity of around $0.9$ when $n=2000$ (except for target FDR level $0.05$ where all catching sets are of size 1). Adding the intermediate layer (layer 1) could further refine catching sets to average size of $1.7$ and purity of $0.95$. However, the power of KeLP is significantly lower especially under lower target FDR levels.
In contrast, by simultaneously selecting important groups and refining the catching sets, the FVG filter return catching sets that are smaller and purer without significant power loss in identifying important genetic variants. As sample size $n$ grows, FVG filter maintains catching sets of small size (around $1.1$) and high purity (higher than $0.97$) for target FDR levels $0.05$, $0.1$, $0.15$, $0.20$. With minor power loss with respect to group knockoff filter, our FVG filter can locate the real important variants from important groups in informative catching sets most of the time. On the contrary, variant-level (layer 0) inferences of KnockoffZoom and KeLP still rely on the same variant-level knockoff of model-X knockoff filter. With low power at layer 0, a non-negligible fraction of catching sets are of size larger than $1$, leading to both large average size and lower average purity. Such an obstacle is elaborately circumvented by our FVG filter which uses group knockoffs to ease the identification of signals and lasso regression to refine the catching sets.
3.4 Multiple Signals in the Same Group
In the simulation setting of Section 3.1, there is only one signal within each nonnull group in most of time. As a result, heuristic identifying the contributing variants in nonnull groups are not that difficult. For example, one can always perform heuristic selection of contributing variants from variant groups identified by the group knockoff filter. However, such an identification task becomes nontrivial when there are multiple genetic signals in one variant group. To demonstrate the performance of the FVG filter in prioritizing important variants under such scenarios, we conduct more experiments with two settings of signal distributions, including:
  • Setting A: $k=15$ selected coefficients ${\beta}_{j}$’s are nonzero and follow $N(0,0.6^{2})$, where there is a randomly selected variant group containing 4 signals, two randomly selected groups containing 2 signals and seven randomly selected groups containing 1 signal;
  • Setting B: $k=20$ selected coefficients ${\beta}_{j}$’s are nonzero and follow $N(0,0.6^{2})$, where there are ten randomly selected groups containing 2 signals.
Under both settings, we simulate 1000 datasets of sample size $n=2000$. We compare our FVG filter with two heuristic approaches.
  • Top-1 signal: Within each variant group selected by the group knockoff filter, we construct the catching set by heuristically including only the variant with the largest $|W_{j}|$’s.
  • SuSiE: Within each variant group selected by the group knockoff filter, we implement the SuSiE method (Wang2020) and construct the catching set as the $95\%$ credible set.
Performance of our FVG filter and two heuristic approaches are presented in Figure 4. We find that the FVG filter still achieves a balance between the informativeness of catching sets and power. On one hand, selecting only the top-1 signal within identified variant groups provides the most informative catching sets (with size 1 and purity 1). However, doing so may miss many signals, especially under setting B. On the other hand, using SuSiE within selected variant groups usually returns too large catching sets. In contrast, our FVG filter not only returns informative catching sets (average size <1.5<1.5 and average purity $\approx 0.9$) but also has similar power as SuSiE. That is to say, among all genetic variants selected by SuSiE that are not top signals in the groups, our FVG filter adaptively selects those contributing variants without proxy ones.
Remark 2.
Although all the simulations are conducted under the linear model, the FDR control of our FVG filter still stands for other conditional distribution $Y|\textbf{X}$. The reason is that our FVG filter is developed under the similar framework of model-X knockoff, where no assumption is imposed on $Y|\textbf{X}$. Thus, even $Y|\textbf{X}$ follows a logistic regression model, using the lasso feature importance scores would not violate our FVG filter’s FDR control. However, using a model that accurately depicts $Y|\textbf{X}$ to compute feature importance scores would further improve the power.
Refer to captionRefer to captionRefer to caption
(a) Setting A.
Refer to captionRefer to captionRefer to caption
(b) Setting B.
Figure 4: Average power, sizes and purity of catching sets obtained by FVG filter over 1000 simulated genetic datasets of different settings in comparison to two heuristic approaches.